# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

### v1

@@ A.6 Developed sensing applications
To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Fig. 5. These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch. The data collection is approved with ethics approval and involves 40 subjects. The corresponding target sensing datasets are shown in Tab. 2 with detailed specification.
@@ A.7 Hardware setup and computation environment
We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Tab. 7). For example, Fig. 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4. In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0. We also employ servers equipped with NVIDIA GeForce RTX 3090 GPU to build our cloud environment.
@@ A.8 Training details
In the Repair stage, we fine-tune the generative transfer module (c.f. Sec. 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95. We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5. We also use standard early stopping with a patience of 20. We retain the training parameters of the baselines in Tab. 1 as specified in their original implementations.
@@ B.1 Default reshaping
To align with source model input shape, we develop a default reshaping to transform sensor data shape. It uses bilinear interpolation [84] (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels [85], placed between source models and sensor data.
@@ B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques [46, 47] typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
@@ B.3 Setups for efficient search insights
To gain design insights of efficient search for LWS control (c.f. Sec. 5), we obverse a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Tab. 3. Fig. 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner. The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective). This indicates a strong correlation between the After state’s S-score and both Before and Anchor values. These insights lead to the observation for designing the repair rate growth model.
@@ B.4 Observation for designing repair rate growth model
Our key idea is to estimate the repaired S-score (i.e., repair value) of each layer before initiating SRR processing, allowing us to decide whether to bypass unnecessary repairs. To empirically examine this, we first define a repair rate function to capture this correlation as $rate_{n}=(After-Before)/Anchor=(V_{ij}^{n}-Score(Pro(ft_{ij}^{n}),Y_{t}))/Score(Pro(fs_{ij}^{n}),Y_{s})$, where $Y_{s}$ denotes the source labels, and we normalize $rate_{n}\in[0,1]$. We then repeat the process multiple times to observe the changes in $rate_{n}$. Fig. 6(c) shows that the repair rate demonstrates exponential growth. Thus, we design the repair rate growth model based on an exponential growth model (c.f. Sec. 5.2).
@@ C.1 Removal performance
We evaluate how channel removal enhances the average S-score of selected layers. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting, and enable channel removal with a default search depth of 3. We employ L2 norm-based (i.e., L2-based) model pruning [47, 50] as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Fig. 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.2 Impact of search depth
@@
In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models. Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy. To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Tab. 2) in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9. Fig. 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieves low FLOPs. As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs. Although both Single-D3 and Multi-D3 do not produce the lowest FLOPs on average, they outperform others and achieve the highest average model accuracy of 70.52% and 70.69%, respectively, across the three settings. Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
C.3 Regression task performance
@@
We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Tab. 2 collected by the blood pressure monitoring application. We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting. To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing. For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg [86]. We employ a random forest (RF) based method [87] as the oracle baseline. Besides, we train a Conv4 backbone in Tab. 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines. Fig. 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard. In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks. These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
C.4 On-device performance
We now evaluate the on-device performance of deploying XTransfer’s output models on 3 commercial edge devices (Tab. 7), focusing on resource overhead metrics such as FLOPs, parameters, latency per call, and peak memory footprint. For each target dataset, we test both types of output models created by our single-source (Single) and multi-source (Multi) methods. We also use ResNet18 and Multi-ResNet18 (Tab. 3) as baselines. As shown in Tab. 8, all output models achieve significant reductions in model resource overhead, e.g., FLOPs is reduced by 2.4 to 5.7$\times$ for Single models and by 10.1 to 16.5$\times$ for Multi models, compared to the baselines. These recombined models also demonstrate substantial improvement in on-device performance. Specifically, in the Single method, latency is reduced by 1.4 to 4.1 $\times$ and peak memory footprint decreases by 1.8 to 3.9$\times$. In the Multi method, the reductions are even more pronounced, with latency dropping by 7.4 to 21$\times$ and memory footprint by 2.2 to 8.4$\times$, compared to the baselines. These results prove that XTransfer successfully achieves streamlined edge deployment.

### v2

@@ A.6 Developed sensing applications
To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Figure 5. These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch. The data collection is approved with ethics approval and involves 40 subjects. The corresponding target sensing datasets are shown in Table 2 with detailed specifications.
@@ A.7 Hardware setup and computation environment
We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Table 7). For example, Figure 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4. In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0. We also employ servers equipped with NVIDIA GeForce RTX 3090 GPU to build our cloud environment.
@@ A.8 Training details
In the Repair stage, we fine-tune the generative transfer module (see Section 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95. We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5. We also use standard early stopping with a patience of 20. We retain the training parameters of the baselines in Table 1 as specified in their original implementations.
@@ B.1 Default reshaping
To align with source model input shape, we develop a default reshaping to transform sensor data shape. It uses bilinear interpolation (Mastyło, 2013) (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels (Szegedy et al., 2014), placed between source models and sensor data.
@@ B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques (Shen et al., 2022; Frankle & Carbin, 2019) typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
@@ B.3 Setups for efficient search insights
To gain design insights of efficient search for LWS control (see Section 5), we observe a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Table 3. Figure 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner. The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective). This indicates a strong correlation between the After state’s S-score and both Before and Anchor values. These insights lead to the observation for designing the repair rate growth model.
@@ B.4 Observation for designing repair rate growth model
Our key idea is to estimate the repaired S-score (i.e., repair value) of each layer before initiating SRR processing, allowing us to decide whether to bypass unnecessary repairs. To empirically examine this, we first define a repair rate function to capture this correlation as $rate_{n}=(After-Before)/Anchor=(V_{ij}^{n}-Score(Pro(ft_{ij}^{n}),Y_{t}))/Score(Pro(fs_{ij}^{n}),Y_{s})$, where $Y_{s}$ denotes the source labels, and we normalize $rate_{n}\in[0,1]$. We then repeat the process multiple times to observe the changes in $rate_{n}$. Figure 6(c) shows that the repair rate demonstrates exponential growth. Thus, we design the repair rate growth model based on an exponential growth model (see Section 5.2).
@@ Appendix C Additional experiment results
C.1 Feature space alignment performance
In the process of feature space alignment (Section 4.1), we use a standard Hungarian algorithm based on default distance metrics to select and pair classes. To verify its effectiveness, we keep using the setup of HHAR, the standard pre-trained ResNet18 on miniImageNet, and the standard PCA under 5-shot settings to examine how our repair loss performs with/without the feature space alignment, especially in the initial layer repairing. Repairing the initial layer is critical, as it often introduces a large, disorganized feature space and significantly affects the repair of subsequent layers. In Table 8, the results show that with the alignment, the repair loss decreases considerably faster by 1.87 times on average, compared to the case without alignment. This indicates that the design of feature space alignment notably contributes to the repairing process.
Epoch (#) #1 #20 #40 #60 #80 #100
With Alignment 0.486 0.355 0.350 0.339 0.333 0.331
Without Alignment 0.476 0.406 0.393 0.383 0.378 0.387
Table 8: Feature space alignment performance across training epochs.
C.2 Impact of PCA variations
Since our method uses the repair learning mechanism that relies on backpropagation (BP) through the reduced-orthogonal feature space, standard PCA can project data into a linear subspace and maintain a closed-form, differentiable structure, which enables smooth integration with BP in our end-to-end learning pipeline. Compared to advanced choices such as kernel PCA, we find that it introduces a non-linear mapping into a high-dimensional feature space using a kernel function, and its projections are generally not differentiable with respect to input features. This makes it incompatible with our pipeline. To test whether linear PCA variations affect the overall performance of XTransfer, we compare standard PCA against sparse PCA using the setup of HHAR based on the ResNet18 under 3 few-shot settings with a default search depth of 3. As shown in Table 9 (#2 denotes 2-component setting), sparse PCA reduces learning time slightly by 11.6% on average, but suffers from a notable accuracy drop by 4.4% on average. These results support the choice of standard PCA as a balanced and practical solution.
PCA Variation 3-shot 5-shot 10-shot
Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
Standard PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
Sparse PCA-#2 58.52 3.8 64.67 5.3 75.11 8.8
Table 9: Impact of PCA variations on accuracy and training time.
C.3 Impact of PCA components
We then evaluate the impact of PCA components on the model accuracy and time efficiency of XTransfer. Since the feature space with different dimensionalities adjusted by PCA components may impact our method performance, we test 3 different levels of PCA components, including PCA-#2 (top two components), PCA-80% (selected components covering 80% of the feature variance), and PCA-Max (maximum components available). We continue to use the same setup as above with the standard PCA. The results in Table 10 show that as the number of PCA components increases, the training time rises notably. In fact, PCA-Max results in an average accuracy drop of 4.5% compared to PCA-80%. By contrast, PCA-#2 achieves the best accuracy-to-time ratio across all n-shot settings, with competitive accuracy and the lowest training time, hence we use PCA-#2 as the default throughout the evaluation.
PCA Components 3-shot 5-shot 10-shot
Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
PCA-80% 63.61 5.2 74.72 7.6 80.00 12.9
PCA-Max 60.00 7.2 75.56 8.1 70.28 16.9
Table 10: Impact of PCA components on accuracy and training time.
@@ C.4 Removal performance
We next evaluate how channel removal enhances the average S-score of selected layers. We continue to use the same setup as above under 5-shot settings. We employ L2 norm-based (i.e., L2-based) model pruning (Frankle & Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Figure 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.5 Impact of search depth
@@
In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models. Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy. To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Table 2) in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9. Figure 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieve low FLOPs. As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs. Although both Single-D3 and Multi-D3 do not produce the lowest FLOPs on average, they outperform others and achieve the highest average model accuracy of 70.52% and 70.69%, respectively, across the three settings. Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
C.6 Impact of model backbones
The current implementation of XTransfer uses homogeneous pre-trained source models. Technically, the layer-wise repairing and recombination techniques impose no fundamental restrictions on the choice of model backbone and work with various architectural variants. To verify this, we evaluate the trade-off between accuracy and computational cost using different source backbones. We continue to use HHAR, and test Multi-Conv4, Multi-ResNet10, and Multi-ResNet18 in a 5-shot, multi-source setting shown in Table 3. The results in Table 11 demonstrate that the average accuracy declined, transitioning from Multi-Conv4 to Multi-ResNet10, and then increasing with Multi-ResNet18. Interestingly, Multi-Conv4, despite having fewer parameters, outperforms Multi-ResNet10, which may be attributed to differences in model quality or distributional divergence between source and target. Multi-ResNet18, though with higher FLOPs, achieves the best accuracy. It suggests that larger models trained on large-scale cross-modality datasets exhibit superior latent features, which notably contribute to such model transfer. Our techniques such as the connectors are designed to flexibly handle the layers with different shapes/structures, and can be extended to heterogeneous model backbones. In addition, our method focuses on latent feature alignment rather than specific feature extraction. It remains fully compatible with advanced source models (e.g., with temporal encoders), which can be seamlessly integrated into our pipelines with minimal or no modifications.
Backbone Accuracy (%) FLOPs (G)
Multi-Conv4 65.56 0.01
Multi-ResNet10 63.61 0.14
Multi-ResNet18 78.44 2.19
Table 11: Performance of different source model backbones in terms of accuracy and FLOPs.
C.7 Training efficiency
To test the practical training efficiency in the cloud, we further evaluate training time across the selected baselines. We use the same experimental setups as in Table 6, Section 6.3. For MAML (Chen et al., 2019) and SemiCMT (Chen et al., 2024), which require model re-training from pre-trained source models to adapt to the target modality classes, training time includes both re-training and fine-tuning. For DAPN (Zhao et al., 2021), it reflects model adaptation process on few target data. The results in Table 12 show that both Our-Single and Our-Multi achieve significant training time reductions compared to MAML. Our-Single also matches or exceeds DAPN and SemiCMT in training efficiency. Notably, even when using up to 5 source models, Our-Multi remains practical training efficiency, requiring only 2.05 times more training time on average compared to Our-Single across all datasets and shot settings. This further supports the practicability and scalability of our method.
HHAR (mins) WESAD (mins) Gesture (mins) Writing (mins) Emotion (mins) ChestX (mins)
DAPN (Zhao et al., 2021) 2.38 4.62 11.78 1.49 2.54 6.78 3.36 6.12 15.72 4.12 7.71 20.97 2.99 7.34 18.86 2.14 4.57 12.04
MAML (Chen et al., 2019) 124.28 125.66 183.20 71.71 77.07 115.87 118.79 129.79 191.95 165.06 129.27 194.36 94.81 98.40 153.20 66.20 72.17 106.97
SemiCMT (Chen et al., 2024) 3.02 8.27 10.23 1.82 2.48 5.63 4.81 9.42 12.83 10.10 13.92 18.75 3.39 7.32 11.48 2.84 5.28 8.33
Our-Single 3.98 5.57 7.95 2.12 2.99 4.33 5.64 7.80 11.24 7.94 10.54 14.69 5.58 7.43 9.88 3.65 4.75 6.84
Our-Multi 9.25 15.33 15.25 4.40 5.38 7.21 12.49 15.12 24.37 15.35 17.36 32.38 11.72 13.39 19.75 8.41 9.33 13.61
Table 12: Comparison in training time of each method.
C.8 Regression task performance
@@
We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Table 2 collected by the blood pressure monitoring application. We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting. To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing. For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg (Stergiou et al., 2018). We employ a random forest (RF) based method (Shi et al., 2023) as the oracle baseline. Besides, we train a Conv4 backbone in Table 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines. Figure 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard. In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks. These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
C.9 On-device performance
[…]